# smile: Single-turn to Multi-turn Inclusive Language Expansion via ChatGPT for Mental Health Support

### v2

@@ Appendix H Implementation Details
Table 10: Parameters of parameter-efficient fine-tuning.
Dialogue Generation
@@
We use ChatGLM2-6B as the baseline model, and all hyperparameters during generation are set to their default values from the official repository [^6] https://huggingface.co/THUDM/chatglm2-6b. After tuning ChatGLM2-6B with LoRA, we set the maximum generation length to 2040, temperature to 0.8, and adopt nucleus sampling with $p=0.8$.
Appendix I Prompt Details
Prompt A is directly utilized from an online GitHub repository, while prompt B is designed by our professional counselors.
Prompt A [^7] https://github.com/PlexPt/awesome-chatgpt-prompts-zh
我想让你担任心理医生。我将为您提供一个寻求指导和建议的人,以管理他们的情绪、压力、焦虑和其他心理健康问题。您应该利用您的认知行为疗法、冥想技巧、正念练习和其他治疗方法的知识来制定个人可以实施的策略,以改善他们的整体健康状况。 (English translation: I want you to take on the role of a psychologist. I will provide you with someone seeking guidance and advice to manage their emotions, stress, anxiety, and other mental health issues. You should use your knowledge of cognitive-behavioral therapy, meditation techniques, mindfulness practices, and other therapeutic methods to formulate strategies that individuals can implement to improve their overall health.)
Prompt B
现在你扮演一位专业的心理咨询师,你具备丰富的心理学和心理健康知识。你擅长运用多种心理咨询技巧,例如认知行为疗法原则、动机访谈技巧和解决问题导向的短期疗法。以温暖亲切的语气,展现出共情和对来访者感受的深刻理解。以自然的方式与来访者进行对话,避免过长或过短的回应,确保回应流畅且类似人类的对话。提供深层次的指导和洞察,使用具体的心理概念和例子帮助来访者更深入地探索思想和感受。避免教导式的回应,更注重共情和尊重来访者的感受。根据来访者的反馈调整回应,确保回应贴合来访者的情境和需求。请为以下的对话生成一个回复。 (English translation: Now, you are playing the role of a professional psychological counselor with extensive knowledge of psychology and mental health. You excel in applying various counseling techniques, such as principles of cognitive-behavioral therapy, motivational interviewing skills, and solution-focused short-term therapy. With a warm and friendly tone, demonstrate empathy and a profound understanding of the visitor’s feelings. Engage in a natural conversation with the visitor, avoiding overly long or short responses to ensure a smooth and human-like dialogue. Provide deep guidance and insights, using specific psychological concepts and examples to help the visitor explore thoughts and feelings more deeply. Avoid instructive responses and focus more on empathy and respecting the visitor’s feelings. Adjust responses based on visitor feedback to ensure they align with the visitor’s context and needs. Please generate a response for the following dialogue.)
Appendix J Instructions for Human Evaluation
@@ Appendix I Prompt Details
To maintain the fairness of model evaluation, three responses randomly appear in a different order every time. Furthermore, three professional psychologists are willing to evaluate the response quality, ensuring the quality of human evaluation.
Pairwise human evaluation results, comparing the baseline model to ground truth, are reported, as illustrated in Figure 12. Results show that the baseline model, without training with SmileChat, lags significantly behind compared to the ground truth.
Figure 12: Pairwise human evaluation results, comparing the baseline model to ground truth, are reported. We present the win and lose rates of each compared pair in 100 randomly sampled real-life dialogue sessions. Fleiss’ kappa (Fleiss et al., 1981) is used to measure the inter-rater agreement, and the value falls within moderate agreement with $0.5\leq\kappa\leq 0.6$.
Appendix K Definitions of Dialogue Topics
1. Mate Selection: Do not know how to enter into an intimate relationship, have fear or aversion towards intimate relationships, are unwilling to date, or encounter some difficulties in choosing a spouse.
2. Love Issues: Problems encountered by individuals in a romantic state during the process of love, including both long-distance and non-long-distance relationships.
3. Post-love Issues: Interpersonal relationship management with ex-partners after the end of an intimate relationship.
4. Marriage Issues: Limit to relationship problems between spouses. Marriage issues refer to a series of problems that arise between a husband and wife after marriage. They generally include extramarital affairs, emotional infidelity, domestic violence, personality defects, personality disorders, mother-in-law and daughter-in-law relationships, dull marriage, interpersonal communication, cultural differences, sexual life, and problems in marriages caused by practices like polygamy or polyandry.
5. Concepts of Sex Distress: Sexual concept distress refers to the distress present in understanding and views regarding sexual physiology, psychology, behavior, morality, and civilization.
6. Sexual Preference Distress: Sexual preference distress refers to the psychological distress caused by paraphilias such as fetishism, transvestic fetishism, exhibitionism, friction fetishism, voyeurism, zoophilia, pedophilia, sadomasochism, and necrophilia.
7. Gender Identity/Cognition: It refers to the deeply felt gender within a person, based on personal experience, which may align with the gender assigned at birth (i.e., cisgender) or differ from it (i.e., transgender).
8. Sexual Orientation: Refers to the gender(s) a person is attracted to (including the opposite sex, same-sex, no specific sex, both sexes).
9. Family Conflict: Conflicts, disputes, and communication issues among family members, including those originating from both the nuclear family and the extended family.
10. Child Education: Literally meaning the education of children.
11. Domestic Violence: Physically, mentally, or otherwise harmful actions carried out between family members, including acts such as physical assault, binding, harming, and restricting personal freedom, as well as consistent verbal abuse, intimidation, and similar behaviors.
12. Sexual Harassment: Sexual harassment refers to using sexually suggestive language or actions towards the target of harassment, coercing the victim to comply and making them uncomfortable.
13. Sexual Assault: Sexual assault involves various forms of unwanted sexual contact and coerced sexual activities, including rape, forced kissing, sexual harassment, sexual abuse, exhibitionism, voyeurism, and such actions might also be considered sexual assault in legal precedents.
14. Bullying: Typically refers to the harassment and oppression between individuals with unequal power, which has long existed in society. This includes physical or verbal attacks, resistance and exclusion in interpersonal interactions. It can also involve talking about someone in a sexual harassment-like manner or mocking and making comments about their body parts, or it could be insults and sarcasm driven by personal reasons like jealousy.
15. Loss: The loss of significant others or pets.
16. Setbacks: Inevitable emotional reactions resulting from the obstruction of purposeful actions by individuals. It can cause substantial harm and manifest as disappointment, pain, distress, and unease.
17. Political Violence: Rooted in class domination and relying on the state as its entity, political coercion is the force exerted by the state, relying on various means of coercion such as the military, police, courts, and prisons.
18. Secondary Trauma: After witnessing a large number of cruel and destructive scenes, the degree of psychological and emotional distress exceeds the tolerance limit of some individuals, leading to various abnormal psychological phenomena. These phenomena are often a result of sympathy and empathy for survivors and their trauma, causing severe mental and emotional distress and even psychological breakdown. The main symptoms of secondary trauma include loss of appetite, easy fatigue, decreased physical energy, sleep disturbances (difficulty falling asleep, easy awakening), nightmares, irritability or anger, easy startle response, lack of concentration, feeling numb, fearful, and hopeless about one’s own experiences, accompanied by trauma reactions and interpersonal conflicts.
19. Trauma from Major Life Events: Psychological shadows caused by significant life events other than loss.
20. Psychological Counseling Trauma: Harm caused by therapists during psychological counseling.
21. Health Issues: Life and psychological distress caused by diseases such as heart disease, thyroid nodules, polycystic ovary syndrome, et al.
22. Psychosomatic Symptoms: Including palpitations (a subjective feeling of discomfort with heartbeats), sleep problems (such as insomnia/nightmares), which, if caused by sleep issues leading to other problems, are categorized as sleep problems; eating issues (such as loss of appetite/emotional eating); memory problems; stomach pain; dizziness; fainting; difficulty breathing; physical weakness/tiredness without apparent reason; unexplained bodily pains et al.
23. School/Workplace Adaptation: Literal meaning.
24. Role Transition Adaptation: For example, new mothers/fathers, new wives/husbands, retirees.
25. Cultural Adaptation: Adaptive issues encountered during the ongoing direct contact between individuals from different cultural groups in the process of changes occurring in one or both parties due to their original cultural backgrounds.
26. Self-exploration and Growth: Self-exploration and growth in the course of life development (adolescence, early adulthood, middle age, late adulthood).
27. Personality Trait Exploration: Traits, formation, origins, influences, et al., of personality and character.
28. Exploration of Negative Self-evaluation: Not knowing how to love oneself, feelings of inferiority, low self-esteem, self-denial, self-doubt, internal conflicts, self-contradictions, fearing being different, sensitivity, lack of security, feeling inadequate, et al.
29. Exploration of Life Meaning: Sense of meaninglessness, existential emptiness.
30. Emotion Regulation/Control Methods Consultation: Literal meaning.
31. Depression: An emotion characterized by "low mood, slow thinking, reduced speech and movement."
32. Anxiety: A restless emotion arising from excessive worry about the safety of loved ones, one’s own life, future, destiny, et al.
33. Stress: Stress is a cognitive and behavioral experiential process formed by stressors and stress responses, namely psychological stress.
34. Obsession: Marked by an excessive pursuit of perfection, precision, and a tendency to rationalize conflicts. It involves strong self-control tendencies and self-regulatory behavior to the point of entanglement and nitpicking. Behaviorally, it adheres excessively to rules, regulations, and orders, to the extent that even life details are sought to be procedural and ritualistic, demanding step-by-step compliance.
35. Fear: A strong and repressed emotional state felt deeply by an individual or group in the face of real or imagined danger. Manifestations include heightened nervousness, overwhelming inner fear, inability to concentrate, mental blankness, impaired judgment or self-control, and increased impulsiveness.
36. Dilemma/Decision-Making Difficulty: Faced with numerous choices and concerns, dealing with dilemmas or even multiple predicaments, struggling to make decisions.
37. Impulsiveness/Loss of Control: Often refers to acting recklessly without considering consequences. It involves intense emotions and weak rational control.
38. Interpersonal Communication Skills/Methods Consultation: Inquire about handling interpersonal relationships and more.
39. Interpersonal Conflicts/Disputes: Conflicts, disagreements, dissatisfaction, or communication issues arise during interpersonal interactions. If there are other issues caused by interpersonal relationships, they are also categorized as interpersonal interactions.
40. Social Difficulties/Fears: Characterized by involuntarily feeling nervous and frightened when interacting with others (especially in public settings), leading to confusion, incoherent speech, and even severe fear of social encounters.
41. Learning Efficiency/Methods: Literal meaning.
42. Work Efficiency/Methods: Literal meaning.
43. Job Dissatisfaction: Includes dissatisfaction with salary, benefits, environment, systems, personnel, et al. in a work setting.
44. Learning Dissatisfaction: Includes dissatisfaction with interpersonal relationships, environment, systems, et al. in an educational setting.
45. Occupational Burnout: Refers to the state of physical and mental fatigue and exhaustion experienced by individuals under heavy work pressure.
46. Learning Fatigue: Refers to a phenomenon where students hold a negative attitude towards school courses and academics, accompanied by the following behavioral manifestations: loss of enthusiasm for academic tasks and school activities, displaying a negative state, and showing indifference and estrangement towards classmates and friends.
47. Occupational Stress/Challenges: Various pressures/challenges that arise or form in the workplace, including stress/challenges caused by factors such as heavy workload, difficulties in interpersonal communication, and the impact of changes in the work environment.
48. Academic Pressure: Refers to the mental burden that individuals bear during learning activities. This includes various tense stimuli from the environment experienced during the learning process, as well as measurable and assessable abnormal reactions in the physiological, psychological, and social behavior of students.
49. Dropping Out: Stopping attending school midway or losing the opportunity to attend school.
50. Underemployment/Unemployment: Underemployment refers to the behavior of not finding a job and waiting for job opportunities. Unemployment refers to the situation of a person within a certain age range who is willing and capable of working for remuneration but has yet to find a job.
51. Career Planning: Refers to the ongoing systematic process of planning for one’s career and even life. It includes three elements: career orientation, goal setting, and pathway design.
52. Professional/Career/Occupational Choices/Changes: Literal meaning.
53. Suspected Neuroses and Mental Disorders: Highly suspected of having neuroses and mental disorders, recommended type of consultation.
54. Neuroses and Mental Disorders: It is a general term for a group of mental disorders, including neurasthenia, obsessive-compulsive disorder, anxiety disorder, phobia, somatic symptom disorder, et al. Patients experience significant distress and impairment in psychological or social functioning, but there is no confirmed organic pathological basis. The course of the illness is mostly prolonged or episodic. Mid- to long-term consultations are needed. Mental disorders refer to a range of conditions that meet the diagnostic criteria of DSM-5 (substance addiction or dependence also included) and require hospital consultation.
55. Self-harm/Suicidal Tendencies: Describing incidents of suicide/self-harm/self-mutilation/contemplating suicide. Triggers crisis alerts and referrals.
56. Harming Others/Killing: Describing incidents of harming others/killing/thinking of killing.
Appendix L Dialogue Examplars
Multi-turn dialogue examples generated with the standard and stand.${}_{\mathrm{DT}}$ methods are illustrated in Figure 13 and 15. Further, a dialogue generated by the smile method is shown in Figure 17.
@@
Figure 15: A dialogue example generated with the stand.${}_{\mathrm{DT}}$ method. (Chinese Version)
Figure 16: A dialogue example generated with the stand.${}_{\mathrm{DT}}$ method. (English Version)

### v3

@@ 6.2 Experimental Setup
Table 4: Hyperparameters of parameter-efficient fine-tuning.
Hyperparameters
@@
Table 4 presents the hyperparameters for developing a dialogue model for mental health support. All hyperparameters during inference time are set to their default values from the official repository. After fine-tuning ChatGLM2-6B with LoRA, we set the temperature to 0.8 and top_p to 0.8.
6.3 Evaluation
METEOR $\uparrow$ BLEU-1 $\uparrow$ BLEU-2 $\uparrow$ BLEU-3 $\uparrow$ Rouge-L $\uparrow$ D-1 $\uparrow$ D-2 $\uparrow$ D-3 $\uparrow$ BERTScore $\uparrow$
Baseline 10.15 5.75 1.95 0.85 7.02 52.95 80.74 90.17 56.69
Fine-tuned 14.18 14.03 5.39 2.76 15.83 82.08 95.74 97.81 59.91
Table 5: Results of automatic evaluation in PsyTest dataset. Except for BERTScore, other metrics are evaluated by Chinese character-level tokenization. For BERTScore, we use the BAAI/bge-m3 model to get text embedding (bgem3).
6.3.1 Automatic Evaluation
Metrics
To conduct automatic evaluation, the evaluation metrics we use consist of BLEU-1/2/3 (Papineni et al., 2002), METEOR (Banerjee and Lavie, 2005), Rouge-L (Lin, 2004), Distinct-1/2/3 (D-1/2/3) (Li et al., 2016), and BERTScore(zhang2020bertscore).
Category Total Help-seeker Supporter
# Dialogues 50 - -
# Utterances 3103 1561 1542
Turns per dialogue 31.0 - -
Utterances per dialogue 62.1 31.2 30.8
Avg. length per utterance 32.3 40.2 24.2
# Test cases 1539 - -
Table 6: Data statistics of the dialogue dataset, PsyTest.
Test Set
To better understand and assess the dialogue quality of the SMILECHAT dataset, we propose to utilize real-life multi-turn counseling conversations. We develop an online mental health support platform that enables professional counselors to offer each client a free text-based counseling service, lasting approximately 50 minutes each time. We collect 50 real-life counseling dialogues. To protect user privacy, we ask experts to conduct a data anonymization process, removing information related to user identification (e.g., names and addresses). Then, we split each long dialogue into multiple small sessions, with the last utterance spoken by the counselor. We discard test cases in which the dialogue history is empty, and we name this test set PsyTest, which contains 1539 test cases. Table 6 presents the statistics of the PsyTest dataset.
Results
The results of the automatic evaluation, including nine metrics, are presented in Table 5. Notably, the evaluated dialogues are based on real-world counseling data rather than generated dialogues, which excludes the influence stemming from ChatGPT. All automatic evaluation metrics we use indicate improved performance. Our results show that the model trained with SMILECHAT is effective and practical. Consequently, the automatic evaluation demonstrates that our collected dataset is high-quality.
6.3.2 Human Evaluation
Metrics
We conduct a human evaluation to study the model performance of the dialogue system trained with our proposed dialogue corpus. Initially, we randomly sample 100 test cases from PsyTest, each consisting of a multi-turn dialogue history and a golden response produced by the counselor. Subsequently, we obtain 200 generated responses from the baseline and MeChat models. Three professional counselors are presented with a dialogue history and three randomly shuffled responses (baseline, fine-tuned, ground truth). They are tasked with selecting the optimal response between every two responses for the dialogue history, considering aspects such as professionalism, informativeness, helpfulness, empathy, and safety. The evaluation is conducted based on the ethical principles of psychologists and the code of conduct (Association et al., 2016). For more details about annotation guidelines, refer to Appendix F.
Figure 6: Human evaluation results, including three groups: Golden vs. Baseline, Fine-tuned vs. Baseline, and Fine-tuned vs. Golden. We present the win and loss rates of each compared pair in 100 randomly sampled sessions. Fleiss’ kappa (Fleiss et al., 1981) is used to measure the inter-rater agreement, and all values fall within moderate agreement with $0.5\leq\kappa\leq 0.6$.
Results
We employ majority voting to reach final decisions (win or loss) among three professional counselors. We present human evaluation results, as illustrated in Figure 6, including three groups: Golden vs. Baseline, Fine-tuned vs. Baseline, and Fine-tuned vs. Golden. We present the win and loss rates of each compared pair in 100 randomly sampled sessions. Results show that the baseline model, without training with SMILECHAT, lags significantly behind the setting with golden responses.
As depicted in Figure 6, the model trained with SMILECHAT demonstrates a significant performance improvement compared to the baseline model. Moreover, the responses generated by MeChat surpass the golden response to some extent in 36% of all dialogue sessions. In short, human evaluation also proves that the SMILECHAT dataset is high-quality.
Case Study
We present a case study, as shown in Figure 7. First, the text should not be too long during counseling conversations for the help-seeker to understand. Second, at the beginning of a conversation, a supporter generally will not challenge the help-seeker. Instead, they will provide support first. Third, response 1 is more human-like and natural. Therefore, the first response is better than the second response overall. Therefore, fine-tuning LLMs is indispensable in mental health support.
Figure 7: Case study. Counseling strategies used in the two responses are presented in parentheses. Strategies in green are supportive, while those in red are challenging and should not be used in the early stages of counseling.
7 Conclusion
This paper introduces SMILE, a simple yet effective solution for addressing the data scarcity of multi-turn conversations in mental health support. Through language transformation and diversity analysis, we confirm the feasibility and effectiveness of our approach. The proposed method automatically creates a large-scale, lifelike, diverse, and high-quality dialogue corpus, SMILECHAT, consisting of 55165 dialogues with an average of 5.7 turns. Both automatic and human evaluations using the PsyTest dataset, consisting of 50 real-life anonymized counseling dialogues, demonstrate that SMILECHAT significantly improves dialogue system performance in mental health support. We release multi-turn dialogues (SMILECHAT), a dialogue model (MeChat), and a real-life anonymized test set (PsyTest) to drive progress in the research community.
Limitations
Our study has some limitations that could be addressed in future research.
In this paper, the automatic evaluation does not reasonably reflect the model performance, given the one-to-many problem. In future work, therefore, we will explore a more comprehensive tool kit for evaluating conversational agents in the mental health domain.
Furthermore, there is a limitation in the dialogue turns of the SMILECHAT dataset. Therefore, we will explore other methods to synthesize the dialogue data with more dialogue turns.
Ethics Statement
Our research is reviewed and approved by the Westlake University Institutional Ethics Committee (20211013LZZ001).
Data Sharing
SMILECHAT Based on the data copyright guidelines formulated by PsyQA, we release the multi-turn dialogue corpus publicly available for the research community. If researchers wish to reproduce the multi-turn dialogues using PsyQA, they should sign an agreement with the original data owner. Accordingly, we release our datasets and models for research purposes, thus facilitating further advancement in the academic community.
PsyTest Considering the nature of psychological counseling data, we must cautiously share this dataset. Regarding the rules for releasing data, third-party researchers who require access to the PsyTest dataset must provide us with their valid ID, proof of work, and the reason they are requesting the data (e.g., the research questions). They are required to be affiliated with a non-profit academic or research institution. This includes obtaining the approval of an Institutional Review Board (IRB), having principal investigators working full-time, as well as obtaining written approval from the Institution Office of Research or equivalent office. Additionally, they must sign a Data Nondisclosure Agreement and promise not to share the data with anyone. However, for-profit organizations that want to use this data must sign a license agreement to gain access to the dataset.
Acknowledgements
We thank the anonymous reviewers for their valuable comments. This work was supported by the Research Center for Industries of the Future at Westlake University (Grant No. WU2023C017) and the Key Research and Development Program of Zhejiang Province of China (Grant No. 2021-C03139).
Appendix A Details of Data Cleaning
A.1 Automatic Cleaning
We employ a sequential pipeline for data cleaning to swiftly replace words that are unsuitable to the conversation scenario. {CJK*}UTF8gbsnFor example, both "楼主你" (literally thread starter you) and "楼主" (literally thread starter) should be replaced with "你" (you). However, it is necessary to perform the former replacement to avoid the repetition of "你" and the resulting "你你" (you-you).
A.1.1 Word List for Data Cleaning
To avoid the repetition of {CJK*}UTF8gbsn"你" (you) and the resulting "你你" (you-you) , we suggest to conduct a sequential word replacing pipeline. Figure 8 shows the word list for data cleaning and the corresponding order for automatic cleaning.
Figure 8: Word list for automatic cleaning.
A.2 Manual Cleaning
Due to the specificity and complexity of language, manual cleaning remains an essential part of the process. To prevent virtual dialogue systems from exhibiting overly frequent anthropomorphic behavior, we identify instances of the Chinese word for "hug" ({CJK*}UTF8gbsn抱抱) and manually delete sentence snippets containing this term.
Appendix B Requirements for Dialogue Filtering
B.1 Data Format
We provide the requirements for data format as follows: {CJK*}UTF8gbsn
  • 1.
    The generated conversations do not start with "求助者" or "支持者".
  • 2.
[…]
  • 3.
  • 4.
B.2 Dialogue Turns
Conversations comprising fewer than five turns will be discarded.
Appendix C Method
We present the standard, standardT and SMILE prompts in Figures 9, 10, and 2, respectively.
Figure 9: The standard method used to generate dialogues for mental health support.
Figure 10: The standardT method used to generate dialogues for mental health support.
Appendix D Dialogue Topics Annotation
In this paper, to label the dialogue topics of generated dialogues, the hyperparameters of Qwen1.5-110B-Chat we used are set to the officially recommended default values, where temperature $=0.7$ and top_p $=0.8$. Figure 11 shows the prompting template of dialogue topics annotation.
Figure 11: Prompting template of dialogue topics annotation, where the content in bold is a placeholder.
Appendix E System Prompt Details
We present the system prompt in Figure 12.
Figure 12: System prompt for fine-tuning.
Appendix F Instructions for Human Evaluation
@@
In order to maintain the fairness of model evaluation, three responses randomly appear in a different order every time. Furthermore, three professional psychologists are willing to evaluate the response quality, ensuring the quality of human evaluation.
Appendix G Definition of Dialogue Topics
We present the definition of dialogue topics, as shown in Figures 14 and 15. For English version, refer to Figures 16 and 17.
Figure 14: Dialogue Topics (Chinese version, Part 1).
Figure 15: Dialogue Topics (Chinese version, Part 2).
Figure 16: Dialogue Topics (English version, Part 1).
Figure 17: Dialogue Topics (English version, Part 2).
Appendix H Dialogue Example
Multi-turn dialogue examples generated with the standard and standardT methods are illustrated in Figures 18 and 20. Further, a dialogue generated by the SMILE method is shown in Figure 22. For English version, refer to Figures 19, 21, and 23.
@@ Appendix H Dialogue Example
Figure 20: A dialogue example generated with the standardT method. (Chinese Version)
Figure 21: A dialogue example generated with the standardT method. (English Version)